Eric bikes 12 mph with no wind. Against the wind she bikes 8 miles in the same time that it takes to bike 14miles with the wind. Find the speed of the wind.
Answer by venugopalramana(3286)

Eric bikes 12 mph with no wind. Against the wind she bikes 8 miles in the same time that it takes to bike 14miles with the wind. Find the speed of the wind.
You may edit the question. Maybe convert formulae
speed with no wind=12 mph
wind speed =w
speed against wind =12-w mph
speed with wind=12+w mph
time=t
distance against wind =t(12-w)=8.......................1
distance with wind=t(w+12)=14...................2
eqn.2/eqn.1=(w+12)/(12-w)=14/8=7/4
4(w+12)=7(12-w)
4w+48=84-7w
7w+4w=84-48=36
11w=36
w=36/11=3.27 mph